OpenStudy (anonymous):

triangle ABC ~ DEF a=8 d=22 and b=4 find e

OpenStudy (anonymous):

e=11

OpenStudy (anonymous):

a/b=d/e

OpenStudy (anonymous):

I'ma explain it. A/D = 8/22 B/E = 4/x Cross multiply. 22*4 = 88 8x = 8x 88 = 8x 88 divided by 8 = 11 E = 11

OpenStudy (anonymous):

D: No medal for Emmy.

OpenStudy (anonymous):

Waaaaaaaaaaait

OpenStudy (anonymous):

I did that wrongggg

OpenStudy (anonymous):

..No it's right. Ok!